# Fast quantum gates for exchange-only qubits using simultaneous exchange pulses

### v1

@@ II.4 All-to-all connectivity
In case all interactions are independently tunable, the previously discussed pulsing schemes still apply. Furthermore, all rotation axes lying in the $x-z$ plane are possible in a single simultaneous pulse, as depicted in Fig. 5(b). The two vectors $\vec{n}=-(\sqrt{3},0,1)$ and $\vec{m}=(\sqrt{3},0,-1)$ correspond to $J_{23}$ and $J_{13}$, respectively. In particular a single-pulse implementation of a rotation around the $x$-axis requires $J_{12}=(J_{23}+J_{13})/2$. We find a rotation around the vector $(1,0,0)$ for $J_{23}>J_{13}$ and around the vector $(-1,0,0)$ for $J_{23}<J_{13}$. Thus $x$ rotations around an angle $\theta\in(0,\pi)$ can be realized fastest with $J_{13}=0$ and around $\theta\in(\pi,2\pi)$ can be realized fastest with $J_{23}=0$. Moreover, a frequently used gate in quantum algorithms, the Hadamard gate, can be realized using one simultaneous pulse with a gate time comparable to the $Z$ and $X$ gate. We summarize and compare simultaneous single-qubit gates to sequential pulses and linearly connected qubit architectures in Tab. 1. We point out that whereas excluding the identity the average gate time for sequential pulses is 73.420 ns, for simultaneous pulses in linearly connected quantum dots it is 60.010 ns, while for all-to-all connectivity it reduces to 44.136 ns. Here we assumed rectangular exchange pulses with maximal exchange value $J_{\rm max}=100$ MHz each. As the gate time reduces with simultaneous pulsing the number of applied quantum gates can be increased until the qubit loses its coherence.
Analogously to the linear case, we again consider the subsequences in two-qubit gates given in Tab. 2 and find single-pulse implementations. Since all-to-all connectivity gives an additional degree of freedom it is possible to compensate for phases within one pulse. In Fig. 4(c) the single-pulse gate with all-to-all connectivity equivalent to a $\pi/2-\pi/2-\pi/2$ sequence is showcased.
Furthermore, not only can these selected subsequences be achieved when using simultaneous pulsing with all-to-all connected spins, but we can also engineer the relative phase $\phi$ between qubit and leakage states by diagonalizing the time evolution using $\tau=2\pi/\sqrt{J_{12}^{2}+J_{13}^{2}+J_{23}^{2}-J_{13}J_{23}-J_{12}J_{13}-J_{12}J_{23}}$. Then we obtain a time evolution as in Eqs. (6) and (7) where the phase $\phi_{A}=\phi=(J_{12}+J_{13}+J_{23})\tau/2$ can assume any possible value $0\leq\phi<2\pi$. In Fig. 4(b) $\phi$ is calculated depending on the ratios $J_{13}/J_{12}$ and $J_{23}/J_{12}$.
@@ III Conclusions
In this paper we have shown an alternative pulsing scheme to operate exchange-only qubits using two neighboring exchange pulses simultaneously. We have found conditions for faster single-qubit gates with fewer pulses and compared the linear and all-to-all connected arrangement, and found that the latter offers more flexibility and thus even faster gates.
@@ Appendix A Charge noise sweet spots in triple quantum dots
where $\hat{n}_{i}=\hat{n}_{i,\downarrow}+\hat{n}_{i,\uparrow}$ and $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$ we obtain the exchange interactions between neighboring electron spins
@@
Here we defined $\tilde{V}_{12}=(V_{12}-V_{23}-V_{13})/2$, $\tilde{V}_{23}=(-V_{12}+V_{23}-V_{13})/2$ and $\tilde{V}_{13}=(-V_{13}-V_{23}+V_{13})/2$. We find a first order sweet spot for fluctuations $\delta\epsilon_{i}$ in $\epsilon_{i}$ if

### v2

@@ II.3 All-to-all connectivity
In case all interactions are independently tunable, the previously discussed pulsing schemes still apply. Furthermore, all rotation axes lying in the $xz$-plane are possible in a single simultaneous pulse [29], as depicted in Fig. 5(b). The two vectors $\vec{n}=-(\sqrt{3},0,1)$ and $\vec{m}=(\sqrt{3},0,-1)$ correspond to $J_{23}$ and $J_{13}$, respectively. In particular a single-pulse implementation of a rotation around the $x$-axis requires $J_{12}=(J_{23}+J_{13})/2$. We find a rotation around the vector $(1,0,0)$ for $J_{23}>J_{13}$ and around the vector $(-1,0,0)$ for $J_{23}<J_{13}$. Thus $x$ rotations around an angle $\theta\in(0,\pi)$ can be realized fastest with $J_{13}=0$ and around $\theta\in(\pi,2\pi)$ can be realized fastest with $J_{23}=0$. Moreover, a frequently used gate in quantum algorithms, the Hadamard gate, can be realized using one simultaneous pulse with a gate time comparable to the $Z$ and $X$ gate. We summarize and compare simultaneous single-qubit gates to sequential pulses and linearly connected qubit architectures in Tab. 1. We point out that whereas excluding the identity the average gate time for sequential pulses is 73.420 ns, for simultaneous pulses in linearly connected quantum dots it is 60.010 ns, while for all-to-all connectivity it reduces to 44.136 ns. Here we assumed rectangular exchange pulses with maximal exchange value $J_{\rm max}=100$ MHz each. As the gate time reduces with simultaneous pulsing the number of applied quantum gates can be increased until the qubit loses its coherence.
Analogously to the linear case, we again consider the subsequences in two-qubit gates given in Tab. 2 and find single-pulse implementations. Since all-to-all connectivity gives an additional degree of freedom it is possible to compensate for phases within one pulse. In Fig. 4(c) the single-pulse gate with all-to-all connectivity equivalent to a $\pi/2-\pi/2-\pi/2$ sequence is showcased. In Appendix B we show how three-step sequences can be replaced by simultaneous pulses in the Fong-Wandzura CZ gate.
Furthermore, not only can these selected subsequences be achieved when using simultaneous pulsing with all-to-all connected spins, but we can also engineer the relative phase $\phi$ between qubit and leakage states by diagonalizing the time evolution using $\tau=2\pi/\sqrt{J_{12}^{2}+J_{13}^{2}+J_{23}^{2}-J_{13}J_{23}-J_{12}J_{13}-J_{12}J_{23}}$. Then we obtain a time evolution as in Eqs. (6) and (7) where the phase $\phi_{A}=\phi=(J_{12}+J_{13}+J_{23})\tau/2$ can take any possible value $0\leq\phi<2\pi$. In Fig. 4(b) $\phi$ is calculated depending on the ratios $J_{13}/J_{12}$ and $J_{23}/J_{12}$.
@@ III Conclusions
In this paper we have shown an alternative pulsing scheme to operate exchange-only qubits using two neighboring exchange pulses simultaneously. We have found conditions for faster single-qubit gates with fewer pulses and compared the linear and all-to-all connected arrangement, and found that the latter offers more flexibility and thus even faster gates. Increasing the connectivity between spins comes along with challenges such as crosstalk on the control gate electrodes that can be calibrated using virtual gates [30, 31]. Also the number of residual exchange interactions present during idling times increases. On the other hand, the number of control pulses necessary for a specific gate decreases with increased connectivity. This reduces the number of idling times between two pulses.
@@ Appendix A Charge noise sweet spots in triple quantum dots
where $\hat{n}_{i}=\hat{n}_{i,\downarrow}+\hat{n}_{i,\uparrow}$ and $\hat{n}_{i,\sigma}=\hat{c}_{i,\sigma}^{\dagger}\hat{c}_{i,\sigma}$, we obtain the exchange interactions between neighboring electron spins
@@
Here we defined $\tilde{V}_{12}=V_{12}-V_{23}-V_{13}$, $\tilde{V}_{23}=-V_{12}+V_{23}-V_{13}$ and $\tilde{V}_{13}=-V_{12}-V_{23}+V_{13}$. We find a first order sweet spot for fluctuations $\delta\epsilon_{i}$ in $\epsilon_{i}$ if
@@
Appendix B Simultaneous exchange interaction in two-qubit gates
For the Fong-Wandzura CZ gate as in Ref. [17] (see Fig. 7(a)) we explicitly show how subsequences in Tab. 2 can be used to replace three-step sequences. We show the original CZ gate sequence in Fig. 7(a). If we assume the same linear connectivity as in the initial sequence, shown in Fig. 7(e), we can replace the subsequences as shown in Figs. 7(b) and (c) and obtain Fig. 7(d). On the other hand, if we allow for a triangular connectivity [25] as shown in Fig. 7(g), we can replace the subsequences shown in Figs. 7(h)-(j) in the original sequence and obtain the CZ gate sequence in Fig. 7(f). All three sequences Figs. 7(a), (d) and (f) result in the same gate operation. We again note that the sequence was found and optimized for the brick-structured sequential sequence, The original sequence can be executed in 13 time steps, where within each time step the applied exchange interactions commute. For $J_{\rm max}=100$ MHz the gate time is then $424.115$ ns + $13\tau_{\rm idle}$. The alternative implementation in Fig. 7(d) with linear connectivity requires 16 steps and $680.839$ ns + $16\tau_{\rm idle}$, respectively. The sequence in Fig. 7(f) requires 13 steps and 364.523 ns + $13\tau_{\rm idle}$. When counting each single pulse separately, we find (a) 26 pulses in 753.982 ns + $26\tau_{\rm idle}$, (d) 26 steps in 947.875 ns + $26\tau_{\rm idle}$ of which the phase correction takes 330.279 ns, and (f) 20 steps in 568.727 ns + $20\tau_{\rm idle}$.
Figure 7: (a) CZ gate sequence for a linear connectivity as described in the main text, taken from Ref. [17], first introduced in [10]. Here, $B$ is the control and $A$ the target qubit. (b)-(c) Three-step exchange sequences acting on three spins that can be replaced by two simultaneous exchange pulses. (d) CZ gate sequence equivalent to (a) using the resplacements in (b)-(c). (e) Linear connectivity of the spins used in sequences in (a) and (d). (f) CZ gate for the triangular connectivity in (g) obtained via transformation of the sequence in (a) by replacing the sequences in (h)-(j). (g) Triangular connectivity: Qubit $A$ and $B$ are represented as the left and right triangular, respectively. (h)-(j) Three-step sequences acting on three qubits that can be replaced by a single (simultaneous) pulse. (k) Infidelity of the original sequence in (a), the alternative in linear arrangement in (d) and the triangular arrangement in (f). For the Monte Carlo simulation quasi-static Gaussian-distributed charge noise was introduce with standard deviations up to $\sigma_{ij}=0.1J_{ij}$. (j) Mixing infidelity as defined in Eq. (18)
We estimate the performance of all three CZ gate implementations by calculating the fidelity of each sequence in the presence of noise. The EO qubit is defined in a subspace of a larger Hilbert space. For two-qubit gates states outside of the computational space are used as mediators for the quantum operation. Leakage can arise from charge noise during two-qubit gates and reduces the gate fidelity. However, the resulting actual gate operation that acts entirely outside the computational space is not of particular interest, since leakage states do not carry the desired quantum information. We thus define the subspace averaged gate fidelity by projecting onto the relevant qubit space. The fidelity is defined as [32, 33]
$F=\frac{\text{Tr}[MM^{\dagger}]+\left|\text{Tr}[M]\right|^{2}}{d(d+1)},$ (17)
with $M=PU^{\dagger}_{\rm ideal}PU_{\rm actual}P$, where $P$ is the projector on the subspace of interest with dimension $d$. If leakage has occurred, $M$ is no longer unitary and thus both terms in Eq. (17) give non-trivial contributions [33].
For the fidelity estimation we assume quasi-static Gaussian noise for simplicity and perform a Monte Carlo simulation with standard deviations up to $\sigma_{ij}=0.1J_{ij}$ and 1000 runs. The fidelity of the CZ gates are compared in Fig. 7(k). The original sequence is shown in black, the sequence of Fig. 7(d) with a linear arrangement in blue and the sequence for the triangular arrangement in Fig. 7(f) in red. All curves feature a quadratic slope that can be explained by the quasi-static noise approximation using Gaussian-distributed fluctuations. We find that all three sequences perform similarly and only find a slight improvement in the triangular arrangement. This can be explained by the rather similar length of all sequences, since both new variants were transformed from the original sequence in Fig. 7(a) that was optimized for the sequential execution of non-commuting exchange pairs.
To quantify the mixing between qubit and leakage states we utilize the Frobenius norm $\|A\|=\sqrt{\sum_{i,j}|a_{ij}|^{2}}$ of a matrix $A$. For a unitary matrix $U$ we have $\|U\|^{2}=d$. We thus define the mixing infidelity by
$I_{L}=1-\frac{\|P_{Q}U\|^{2}+\|P_{L}U\|^{2}}{d_{Q}+d_{L}},$ (18)
where $P_{Q}$ and $P_{L}$ are the projections onto the qubit and leakage spaces with dimensions $d_{Q}$ and $d_{L}$, respectively. If $U$ does not introduce any leakage, we have $\|P_{Q}U\|^{2}=d_{Q}$ and $\|P_{L}U\|^{2}=d_{L}$, and thus $I_{L}=0$. However, if $U$ contains block-off-diagonal elements, it introduces transitions between the two spaces and $\|P_{Q}U\|^{2}<d_{Q}$ and $\|P_{L}U\|^{2}<d_{L}$, and hence $I_{L}>0$. Here we neglect the distinction between leakage and seepage, which describe the change of population from qubit to leakage space and vice versa. The mixing infidelity for the Monte Carlo simulations is shown in Fig. 7(l). Again we find a quadratic slope of all curves and only a very slight improvement for the alternative sequences compared to the original one.
Although our results only show a slight decrease in the infidelity for a triangular connectivity, we have shown how the total number of pulses reduces from 26 to 20 which also minimizes the number of idling times between two exchange pulses. More simplifications can be made by optimizing the sequence for simultaneous exchange pulses.